Calculate the mole efficiency of NAOH when 4 grams of NaOH is soluble in 90 grams of water.
1
Expert's answer
2020-01-10T02:31:42-0500

n(NaOH) = m/M = 4g/40g/mol = 0.1 mol

n(H2O) = 90g/18g/mol = 5 mol

The mole efficiency of NaOH = n(NaOH)/n(total) = 0.1mol/5.1mol = 0.02